OpenStudy (anonymous):

Resolve vector v = -i + 4j into two component vectors, one parallel to vector i + j , the other parallel to vector i + 3j.

terenzreignz (terenzreignz):

A systems of linear equations problem in disguise :)

terenzreignz (terenzreignz):

Well, in Maths terms, the problem wants you to express -i + 4j as a sum of multiples of (i + j) and (i+3j) So... -i + 4j = a(i + j) + b(i + 3j) Any ideas as to how to find a and b?

OpenStudy (anonymous):

v=(i+j)a + (i+3j)b v=(ai+aj)+(bi+3bj) v=ai+bi+aj+3bj v=(a+b)i+(a+3b)j i+4j=(a+b)i+(a+3b)j gives us the system -1=a+b and 4=a+3b which has the solutions a=-7/5 and b=2/5 thus, v=\[-i+\left( -\frac{ 1 }{ 5 } \right)j\] done.

terenzreignz (terenzreignz):

It was all good up until you said thus, v= \[\large -i + \left(-\frac{1}{5}\right)j\] How do you figure that?

OpenStudy (anonymous):

From equating a and b into the expression v=(a+b)i+(a+3b)j. v=\[\left( -\frac{ 7 }{5 }a+\frac{ 2 }{ 5}b \right)i+\left( -\frac{ 7 }{ 5 }a+\frac{ 2*3 }{5 } b\right)\] v=\[-i+(-\frac{ 1 }{ 5 })j\]

terenzreignz (terenzreignz):

Well, there's your problem, you're not supposed to end up with a vector v that's different from your original vector. I suggest redoing your system of equations :)

OpenStudy (anonymous):

Yes my system solutions are wrong :) i was studying for 4 hours obvious things seemed complicated :p.. b should equal 5/2 not 2/5 ^^

OpenStudy (anonymous):

Thanks for the pointers ;)